# Measuring Hawking Radiation from Black Hole Morsels in Astrophysical Black Hole Mergers

### v1

@@ Abstract
We show that it is possible to observe the Hawking radiation emitted by small black holes assumed to form in catastrophic astrophysical events such as black hole mergers. Gamma ray bursts in the TeV range are unique footprints of these asteroid-mass black hole morsels ejected during the merger. The time delay of the gamma ray bursts from the gravitational wave event is correlated to the mass distribution of the morsels. The integrated mass of the morsels allowed by the unaccounted merger mass leads to a Hawking induced radiation in photons that is above the sensitivity of atmospheric Cherenkov telescopes such as HESS, LHAASO and HAWC.
Black Holes (BHs) are extreme objects predicted by Einstein’s theory of general relativity. Their presence in the Universe has been detected indirectly via their gravitational interactions. We now know that super-massive Black Holes (SMBHs) are present in the centre of most galaxies [1, 2], with masses between $10^{5}$ and several billion solar masses. They reveal their presence by affecting the rotation speed of nearby stars [3, 4] or producing active galactic nuclei. Recently, the first direct images of the region near a SMBH have been taken [5]. Stellar mass BHs are produced at the end of heavy star lifetimes, and they are at the origin of gravitational wave signals observed at LIGO/VIRGO interferometers [6]. A distinctive signal from BHs may be due, theoretically, to the emission of particles via the evaporation mechanism introduced by S.Hawking [7]. However, since their effective temperature is inversely proportional to their mass, astrophysical BHs heavier than a few solar masses are stable because they are colder than the cosmic microwave background. Smaller BHs, however, emit radiation that can be potentially observed. These asteroid-mass BHs could have been produced in the early Universe [8], and constitute a potential candidate for dark matter if sufficiently long-lived [9, 10, 11, 12]. The Hawking radiation stemming from these primordial BHs is constrained by measuring the diffuse gamma ray background [13].
In this letter, we explore the observational consequences of the production of a large number of small BH morsels during a catastrophic event such as the merger of two astrophysical BHs. As we shall show, the Hawking radiation stemming from these BH morsels gives rise to gamma ray bursts (GRBs) possessing a distinctive fingerprint.
Although BHs lighter than a few solar masses are not expected to be produced via gravitational collapse of over-dense stellar matter, one can still envision that, under certain strong non-linearity in the gravitational field within general relativity or beyond, small BHs may form. For an example, primordial BHs are expected to form via quantum space-time fluctuations enhanced by an inflationary period [14, 15, 16]. Gravitational non-linearity does occur in astrophysical events, as shown in [17, 18, 19, 20] for highly rotating BHs, and in [21, 22, 23] for BH merger dynamics. Specifically, non-linearity appears both in merger simulations [22] and during the successive ringdown phase [21, 23]. BH mergers are especially interesting because they can be detected via gravitational waves. Additionally, the events are expected to be followed by GRBs only in the presence of pre-existing material from their progenitor star [24, 25], from supernova remnants [26, 13, 27] or from active galactic nuclei [28]. The emission of small BH morsels provides a novel mechanism for the emission of high-energy particles, within GRBs characterised by the TeV energy scale.
@@
As we remain agnostic about the details of the BH morsel production, we assume the existence of a distribution of masses that escapes the merger. In other words, when the two BHs merge into a single one, regions of space-time characterised by a strong non-linear gravitational field lead to the formation of BH morsels outside the final horizon. The physics of the Hawking radiation of the BH morsels is encoded in their mass distribution function
$\rho(M_{\rm Bm})\equiv\frac{dn_{\rm Bm}}{dM_{\rm Bm}}\,.$ (1)
@@
$\frac{d^{2}\,N_{p}}{dt\;dE_{p}}=\frac{1}{h}\frac{\Gamma_{p}(E_{p},M_{\rm Bm})}{\exp\frac{E_{p}}{k_{B}T_{\rm Bm}}\pm 1}\,,$ (2)
@@
where $\mu_{t}$ is the mass of the morsels at time $t$ as a function of the initial mass $M_{\text{Bm}}$. The above formula yields the differential primary flux for each particle species $p$. We computed the BH evolution and $J_{p}$ numerically by use of the open-source public code BlackHawk [30, 31], where radiative emission, decays and hadronisation of the primary particles is taken into account by using Pythia [32]. The hadronisation tables, therefore, allow us to reliably generate secondary radiation for primary particle energies between $5$ GeV up to a few TeV. For the numerical results in this letter, we use as template a population of BH morsels with equal masses and non rotating, as angular momentum dissipates faster than the mass of the evaporating BH [33].
For LIGO/VIRGO/KAGRA mergers [34, 35, 36, 37], the BH masses are of stellar order, ranging from a few to several tens of solar masses with associated distance span between 240 Mpc to 3 Gpc. The masses of the initial and final BHs can be indirectly measured from the gravitational wave spectrum, with errors of a few solar masses, depending on the event. For instance, for GW170814 [38], which is the first BH merger observed by all three detectors of LIGO and VIRGO, the initial BH masses have been determined to be $30.5^{+5.7}_{-3.0}$ and $25.3^{+2.8}_{-4.2}$, in units of the solar mass $M_{\odot}$. The final BH mass is $53.2^{+3.2}_{-2.7}\;M_{\odot}$, while the total energy emitted in gravitational waves amounts to $2.7^{+0.4}_{-0.3}\;M_{\odot}$. Taking into account the errors in the parameters, the energy budget of the merger still allows for a total energy of a few solar masses to be emitted in BH morsels. Conservatively, we will take one solar mass as a reference value for the maximal total mass in morsels.
Refer to captionRefer to caption
@@
where we included a geometric factor of $1/2$. In the left panel of Fig. 1 we show the differential fluxes, $E_{\gamma}^{2}dF_{\gamma}/dE_{\gamma}$, for a distribution of BH morsels of same mass, normalised to a total of one $M_{\odot}$. The solid lines provide templates for morsel masses of $2\times 10^{7}$ kg, where the colours indicate different times from the production. As the evaporation lifetime of such mass is about 1 hour, the emission remains constant for timescales of about $500$ sec, emitting the spectrum illustrated by the blue line. Towards the final evaporation time, both the emissivity and the particle energies increase, approaching the ‘explosion’ at the end of the BH lifetime [7]. For the red curve, close to the evaporation time, the primary energies exceed $100$ TeV, reaching the current limit of the BlackHawk code. In fact, the wiggles at high energies correspond to numerical instabilities in the code. We also included the effect of the optical transparency of the intergalactic medium, which cuts off the propagation of photons with energies exceeding $100$ TeV [39], manifesting as a sudden drop in the red curve. For reference, the dotted lines show the primary emission of photons. For different morsel masses, the evaporation time will change drastically as it is proportional to $M_{\text{Bm}}^{3}$. Lighter BHs will, therefore, generate an energetic GRB in a shorter time. Instead, heavier morsels provide a steady signal at the beginning, characterised by reduced particle energies, proportionally to the Hawking temperature $T_{\text{Bm}}\propto M_{\text{Bm}}^{-1}$, while the overall luminosity is reduced. As a reference, we show in dashed green the early curve for a mass of $10^{9}$ kg, corresponding to an evaporation lifetime of about 16 years. Nevertheless, all initial BH morsel masses will generate an energetic GRB with similar characteristics with a time delay from the merger of the order of their evaporation time. In the right panel of Fig. 1 we show the evaporation lifetime as a function of the BH mass, as computed by BlackHawk. Similar fluxes can be obtained for neutrinos.
@@
Multimessenger astrophysics associated to gravitational wave events is part of the LIGO/VIRGO/KAGRA programme, and electromagnetic counterparts are mainly being monitored by Fermi-GBM and Swift-BAT experiments within 30 seconds from all event alerts [40]. Motivated by neutron star mergers, the coverage is between a few keV and MeV. As shown in Fig. 1, we expect very small signals from the BH morsels within this range, well below the current experimental sensitivity of $10^{-7}$ erg sec-1 cm-2 for the flux on Earth. In 2016, however, a short GRB was associated to a BH merger event [26], prompting interest in exploring electromagnetic counterparts of BH mergers [41]. The expected signals are mainly associated with pre-existing material surrounding the BH binary, triggering searches in the optical [42, 43, 44, 45, 46, 47] or keV/MeV [48, 49, 50] ranges. The HESS telescope has followed four BH mergers from the O2 and O3 runs of LIGO/VIRGO [51], covering photon energies between 1 and 10 TeV. However, the observation could only focus on the region of the sky at a time $10^{4}\div 10^{5}$ sec after the BH merger occurred. As no significant point-like source was found in the merger region of the sky, the energy flux from the BH mergers could be constrained to be below $10^{-12}~{}\text{erg}\ \text{sec}^{-1}\ \text{cm}^{-2}$. For morsel masses corresponding to evaporation times within or after the observation window, the HESS data can be used to extract a bound on the total mass that can be emitted as a function of the morsel masses. For each event, we compare the limit on the luminosity to the maximal luminosity of the morsel Hawking radiation within the respective observation window, computed by integrating the source emission between $1$ and $10$ TeV. The results are shown in Fig. 2, where the strongest bounds come from the closest events, GW170814 [38] and GW190728__\__064510. The BH morsel mass ranges stem from the different observation times, where masses smaller than the start of the curves are unconstrained as total evaporation occurs before the HESS observation window. For morsel masses above $15\times 10^{7}$ kg, the energies of the emitted particles fall below the sensitivity range of HESS, and the constraint quickly worsens. This should be taken as a simple estimate, as the limits in [51] were obtained for a $E^{-2}$ power-law spectrum of the photons, which does not exactly match the spectra in Fig. 1.
@@
Figure 2: Estimated upper limit on the total mass $M_{\text{tot}}$ emitted in BH morsels as a function of the morsel mass from the HESS observation of four LIGO/VIRGO events. The curves are extracted from GW170814 (blue), GW190512__\__180714 (red), GW190728__\__064510 (orange) and GW200224 (purple). The upper labels indicate the evaporation time as a function of the BH morsel masses.
@@
As visible signals from BH evaporation always entail photons above the TeV energy, this signal offers a golden opportunity for high energy atmospheric Cherenkov telescopes like HESS, HAWC [52] and LHAASO [53, 54, 55]. The latter recently published a catalogue of multi-TeV sources, reporting the observation of differential fluxes within the range $10^{-13}\div 10^{-11}~{}\text{erg}\ \text{sec}^{-1}\ \text{cm}^{-2}$ [54]. Note that, if the BH morsels have masses above $10^{9}$ kg and evaporation times above several years, BH mergers predating the first LIGO/VIRGO detection would leave behind a relic BH morsel shell and produce a TeV scale GRB uncorrelated to merger catalogues, similar to the signal studied in [56].
@@
Neutrino observatories like ANTARES and IceCube have also monitored BH mergers in search for neutrino counterparts [57]. Both ANTARES [58, 59] and IceCube [60] search for neutrino events within a window of ±$\pm 500$ sec around the BH merger, finding no excess over the expected backgrounds. However, the limits in terms of fluxes are orders of magnitude above the ones for photons, requiring typical source luminosity of $10^{51}$ erg sec-1. As the BH morsels predict similar fluxes in photons and neutrinos, such bounds are not competitive.

### v2

# Looking for Black Hole Morsels in Astrophysical Mergers via Hawking Radiation

@@ Abstract
Gravitational wave observation has provided numerous insights into the merger of astrophysical black holes. In contrast to other violent events (e.g. supernovae), they are, however, not expected to lead to significant emissions of photons and neutrinos. In this paper we discuss a scenario that would lead to characteristic observable gamma ray bursts, which would provide numerous hints to physics beyond General Relativity. Starting from the hypothesis that micro-black holes (called morsels) are formed during the merger process, we show that it is possible to observe their Hawking radiation, which takes the form of gamma ray bursts of a uniquely characteristic form: with energies in the TeV range, their temporal structure is unlike that stemming from any other astrophysical event. Notably, the time delay from the gravitational wave event is correlated to the mass distribution of the morsels. The integrated mass of the morsels, allowed by the unaccounted merger mass, leads to a Hawking radiation in photons that is above the sensitivity of atmospheric Cherenkov telescopes such as HESS, LHAASO and HAWC, and gamma ray space telescopes, such as Fermi-LAT. This renders the hypothesis of morsel creation experimentally testable, and we provide the first concrete bounds on the total mass of morsels formed in specific events.
1 Introduction
Black Holes (BHs) are extreme objects predicted by Einstein’s theory of General Relativity (GR), whose existence has been demonstrated experimentally in numerous different ways: due to their nature, most detections are indirect, i.e. BHs reveal themselves through their gravitational attraction on other (directly observable) objects. In this way, it is by now well established that super-massive Black Holes (SMBHs) of $10^{5}-10^{9}$ solar masses are present in the centre of most galaxies [1, 2], where they significantly affect the rotation speed of nearby stars [3, 4]. Recent advances in large-scale astrophysical experiments, however, have provided new and more direct ways of detecting and measuring BHs. For example, SMBHs can form active galactic nuclei and a first direct image of such a space-time region has been taken in [5]. Furthermore, stellar mass BHs are produced at the end of heavy star lifetimes, and their mergers are at the origin of the gravitational wave signals observed at LIGO/VIRGO interferometers [6].
Further, even more direct ways of directly observing BHs have been theorised within frameworks going beyond GR: concretely, by incorporating quantum effects, a distinctive signal from BHs may be due to the emission of particles via the evaporation mechanism introduced by S. Hawking [7]: based on semi-classical computations, it was argued in this seminal work that BHs have thermodynamic properties and behave like (perfect) black bodies by emitting thermal radiation. A direct observation of this Hawking radiation [^2] Hawking radiation has been experimentally observed in analogue BHs in table-top experiments, see for instance [8, 9, 10]. would give us invaluable insights not only into properties of BHs but also into possible theories of quantum gravity beyond GR. However, it is extremely challenging due to a number of different aspects:
  1. (i)
    The effective temperature is inversely proportional to the BH mass. Hence astrophysical BHs heavier than a few solar masses cannot emit Hawking radiation, because they are colder than the cosmic microwave background. Only the radiation stemming from BHs lighter than half the moon mass, $\sim 3\times 10^{22}$ kg, can be hoped to be observable.
  2. (ii)
    Due to its nature as a black body radiation, it is not directional. Therefore, the flux of Hawking radiation from a (single) distant BH arriving on earth is heavily diluted by the distance squared.
  3. (iii)
    Sufficiently small BHs emit radiation until they completely evaporate: this evaporation process accelerates as the mass diminishes, leading to signals of a burst-like characteristic (with high intensity and high radiation energies) towards the end of a BH lifetime.
These properties suggest that experimentally detectable signals would require very peculiar configurations of BHs, namely small BHs either in close proximity to earth or of sufficient multiplicity to generate a detectable flux. Small BHs weighing like asteroids may have been produced in the early Universe [11] within high density regions. However, these primordial BHs would also have already evaporated during these early cosmological stages and the corresponding Hawking radiation is constrained by measuring the diffuse gamma ray background [12] [^3] Assuming that primordial BHs are long-lived by some mechanisms, they have been discussed as potential candidates for dark matter [13, 14, 15, 16].
While stellar-sized BHs are natural end points of the life-cycle of heavy stars, and even larger ones can be created through consecutive mergers or accretion by matter, possible creation mechanisms for smaller BHs, other than primordial ones, are far more speculative and typically invoke physics beyond GR or beyond the Standard Model of particle physics (see e.g. [17]). In such theories, sufficiently energetic (and potentially catastrophic) astrophysical events (such as the merger of two astrophysical BHs) could produce, as a byproduct, a large number of small BHs. These BH debris, if sufficiently light, evaporate by emitting Hawking radiation and provide an electromagnetic and neutrino signal counterpart to the astrophysical event (such has gravitational waves from BH mergers). The observation of such an event would give us invaluable insights into quantum properties of gravity beyond GR, not only through the measurement of the Hawking radiation, but also by giving hints towards potential creation mechanisms of these small BHs themselves. Therefore, given the obvious inherent theoretical and experiment interest in such a signal, in this paper we shall ask the question: Can such an event be experimentally detected and distinguished from other astrophysical signals?
Concretely, we start from the hypothesis that a significant number of small BHs, can be created during a catastrophic event such as the merger of two astrophysical BHs. [^4] We shall provide additional motivation for this hypothesis in Appendix A, by reviewing fragmentation processes in theories beyond GR. We also provide a very simple back-of-the-envelope estimation for rates and sizes of created BHs by invoking the central dogma [18]. In order to distinguish these from the ‘large’ BH created in the merger (and also primordial BHs), we shall call these black morsels. We thus explore the observational consequences of the production of such morsels during BH mergers: while mergers can be detected via gravitational waves, they are generally not expected to be accompanied by gamma ray bursts (GRB) or high-energy neutrino emission. An electromagnetic counterpart, in fact, is contingent on the presence of pre-existing material from their progenitor star [19, 20], from supernova remnants [21, 12, 22] or from active galactic nuclei [23]. However, as we shall show in this paper, the creation of morsels (and their subsequent evaporation due to Hawking radiation) would lead to a signal of high-energetic photons and neutrinos that carries two marked features:
  • Since the evaporation time of morsels depends on their mass and high energy particles are only emitted towards the end, the onset of the visible gamma ray signal could be significantly delayed with respect to the arrival of the gravitational waves. This delay is correlated to the mass distribution of the morsels. Additional effects may stem from the gravitational field of the merged BH, depending on where the morsels were created.
  • The photon and neutrino energies stemming from morsels typically exceed the TeV scale, unlike astrophysical sources. Moreover, since the wave-length of the Hawking radiation becomes smaller as the morsels approach the point of evaporation, the signal is expected to rise in energy and intensity, before disappearing (when all morsels have evaporated).
We thus argue that the Hawking radiation stemming from these BH morsels gives rise to GRBs possessing a distinctive fingerprint and with energy flux observable at current gamma ray telescopes: this therefore allows for an experimental verification of our hypothesis (i.e. that morsels are produced in BH mergers) at current atmospheric Cherenkov telescopes, such as HESS, LHAASO and HAWC, which are capable of measuring TeV GRBs, and gamma ray space telescopes, such as Fermi-LAT, in the multi-GeV energy range. The uncertainties from gravitational wave detection of BH mergers still allow the emission of a substantial mass (of the order of solar masses) in black morsel debris. Assuming a simple mass distribution of the morsels (motivated by the central dogma [18]), we provide a concrete form of the photon signal that would be experimentally detected on earth. By comparing with actual experimental data, we provide the first upper limits on the total mass emitted in the form of morsels during BH mergers.
While the idea of BH fragmentation during mergers has not, to our knowledge, been directly addressed in the literature, there exists extensive work on post-merger phenomena leading to multimessenger signals. For example, mergers of compact objects could lead to gravitational wave “kicks” [24, 25], accretion disk disruption [26, 27], and transient electromagnetic phenomena [28]. Our scenario differs fundamentally in that it hypothesizes the formation of genuinely distinct compact objects bound by a horizon. We now make this novelty explicit and provide more details on the possible morsel formation in Appendix A.
2 Observable signatures
@@
Since we consider the creation of BH morsels as a (testable) hypothesis, we assume shall assume a distribution of morsel masses that escapes the merger (see also Appendix A.2 for a back-of-the-envelop estimate based on the central dogma). In other words, when the two BHs merge into a single one, regions of space-time characterised by a strong non-linear gravitational field lead to the formation of BH morsels outside the final horizon. The physics of the Hawking radiation of the BH morsels is encoded in their mass distribution function
$\rho(M_{\rm Bm})\equiv\frac{dn_{\rm Bm}}{dM_{\rm Bm}}\,.$ (2.1)
@@
$\frac{d^{2}\,N_{p}}{dt\;dE_{p}}=\frac{1}{h}\frac{\Gamma_{p}(E_{p},M_{\rm Bm})}{\exp\frac{E_{p}}{k_{B}T_{\rm Bm}}\pm 1}\,,$ (2.2)
@@
where $\mu_{t}$ is the mass of the morsels at time $t$ as a function of the initial mass $M_{\text{Bm}}$. The geometric factor of $1/2$ counts the particles emitted away from the merged BH. The above formula yields the differential primary flux for each particle species $p$. We computed the BH evolution and $J_{p}$ numerically by use of the open-source public code BlackHawk [30, 31], where radiative emission, decays and hadronisation of the primary particles is taken into account by using Pythia [32]. The hadronisation tables, therefore, allow us to reliably generate secondary radiation for primary particle energies between $5$ GeV up to a few TeV. For the numerical results in this letter, we use as template a population of BH morsels with equal masses and non rotating, as angular momentum dissipates faster than the mass of the evaporating BH [33].
2.1 Stellar mass mergers
For LIGO/VIRGO/KAGRA mergers [34, 35, 36, 37], the BH masses are of stellar order, ranging from a few to several tens of solar masses, with associated distance span between 240 Mpc to 3 Gpc. The masses of the initial and final BHs can be indirectly measured from the gravitational wave spectrum, with errors of a few solar masses, depending on the event. For instance, for GW170814 [38], which is the first BH merger observed by all three detectors of LIGO and VIRGO, the initial BH masses have been determined to be $30.5^{+5.7}_{-3.0}$ and $25.3^{+2.8}_{-4.2}$, in units of the solar mass $M_{\odot}$. The final BH mass is $53.2^{+3.2}_{-2.7}\;M_{\odot}$, while the total energy emitted in gravitational waves amounts to $2.7^{+0.4}_{-0.3}\;M_{\odot}$. Taking into account the errors in the parameters, the energy budget of the merger still allows for a total energy of a few solar masses to be emitted in BH morsels. Conservatively, we will take one solar mass as a reference value for the maximal total mass in morsels.
@@ 2.1 Stellar mass mergers
In the left panel of Fig. 1 we show the differential fluxes, $E_{\gamma}^{2}dF_{\gamma}/dE_{\gamma}$, for a distribution of BH morsels of same mass, normalised to a total of one $M_{\odot}$. The solid lines provide templates for morsel masses of $2\times 10^{7}$ kg, where the colours indicate different times from the production. As the evaporation lifetime of such mass is about 1 hour, the emission remains constant for timescales of about $500$ sec, emitting the spectrum illustrated by the blue line. Towards the final evaporation time, both the emissivity and the particle energies increase, approaching the ‘explosion’ at the end of the BH lifetime [7]. For the red curve, close to the evaporation time, the primary energies exceed $100$ TeV, reaching the current limit of the BlackHawk code with Pythia secondary modelling. In fact, the wiggles at high energies correspond to numerical instabilities in the code. We also included the effect of the optical transparency of the intergalactic medium, which cuts off the propagation of photons with energies exceeding $100$ TeV [39], manifesting as a sudden drop in the red curve. For reference, the dotted lines show the primary emission of photons. For different morsel masses, the evaporation time will change drastically as it is proportional to $M_{\text{Bm}}^{3}$. Lighter BHs will, therefore, generate an energetic GRB in a shorter time. Instead, heavier morsels provide a steady signal at the beginning, characterised by reduced particle energies, proportionally to the Hawking temperature $T_{\text{Bm}}\propto M_{\text{Bm}}^{-1}$, while the overall luminosity is reduced. As a reference, we show in dashed green the early curve for a mass of $10^{9}$ kg, corresponding to an evaporation lifetime of about 16 years. Nevertheless, all initial BH morsel masses will generate an energetic GRB with similar characteristics with a time delay from the merger of the order of their evaporation time. In the right panel of Fig. 1 we show the evaporation lifetime as a function of the BH mass, as computed by BlackHawk. Similar fluxes can be obtained for neutrinos.
@@
Multimessenger astrophysics associated with gravitational wave events is part of the LIGO/ VIRGO/KAGRA programme, and electromagnetic counterparts are mainly being monitored by Fermi-GBM and Swift-BAT experiments within 30 seconds from all event alerts [40]. Motivated by neutron star mergers, the coverage is between a few keV and MeV. As shown in Fig. 1, we expect very small signals from the BH morsels within this range, well below the current experimental sensitivity of $10^{-7}$ erg sec-1 cm-2 for the flux on Earth. In 2016, however, a short GRB was associated to a BH merger event [21, 41, 42], prompting interest in exploring electromagnetic counterparts of BH mergers [43]. The expected signals are mainly due to pre-existing material surrounding the BH binary system, hence triggering searches in the optical [44, 45, 46, 47, 48, 49] or keV/MeV [50, 51, 52] ranges. The HESS telescope has followed four BH mergers from the O2 and O3 runs of LIGO/VIRGO [53], covering photon energies between 1 and 10 TeV. However, the observation could only focus on the region of the sky at a time $10^{4}\div 10^{5}$ sec after the BH merger occurred. As no significant point-like source was found in the merger region of the sky, the energy flux from the BH mergers could be constrained to be below $10^{-12}~{}\text{erg}\ \text{sec}^{-1}\ \text{cm}^{-2}$. For morsel masses corresponding to evaporation times within or after the observation window, the HESS data can be used to extract a bound on the total mass that can be emitted as a function of the morsel masses. For each event, we compare the limit on the luminosity to the maximal luminosity of the morsel Hawking radiation within the respective observation window, computed by integrating the source emission between $1$ and $10$ TeV. The results are shown in Fig. 2, where the strongest bounds come from the closest events, GW170814 [38] and GW190728__\__064510. The BH morsel mass ranges stem from the different observation times, where masses smaller than the start of the curves are unconstrained as total evaporation occurs before the HESS observation window. For morsel masses above $15\times 10^{7}$ kg, the energies of the emitted particles fall below the sensitivity range of HESS, and the constraint quickly worsens. This should be taken as a simple estimate, as the limits in [53] were obtained for a $E^{-2}$ power-law spectrum of the photons, which does not exactly match the spectra in Fig. 1.
@@
Figure 2: Estimated upper limit on the total mass $M_{\text{tot}}$ emitted in BH morsels as a function of the morsel mass from the HESS observation of four LIGO/VIRGO events [53]. The curves are extracted from GW170814 (blue), GW190512__\__180714 (red), GW190728__\__064510 (orange) and GW200224 (purple). The upper labels indicate the evaporation time as a function of the BH morsel masses. The green band indicates a reinterpretation of the HAWC bound on primordial BHs [54].
@@
As visible signals from BH evaporation always entail photons above the TeV energy, this signal offers a golden opportunity for high energy atmospheric Cherenkov telescopes like HESS, HAWC [55] and LHAASO [56, 57, 58]. The latter recently published a catalogue of multi-TeV sources, reporting the observation of differential fluxes within the range $10^{-13}\div 10^{-11}~{}\text{erg}\ \text{sec}^{-1}\ \text{cm}^{-2}$ [57]. Note that, if the BH morsels have masses above $10^{9}$ kg and evaporation times above several years, BH mergers predating the first LIGO/VIRGO detection would leave behind a relic BH morsel shell and produce a TeV scale GRB uncorrelated to merger catalogues, similar to the signal studied in [59]. Henceforth, a search for BH morsels could follow a similar strategy to searches for local densities of primordial BHs (PBHs). Currently, the strongest bound comes from HAWC [54, 60]. For morsels of same mass, the signal from a morsel distribution at a distance $D$ is equivalent to a single PBH at a distance $D_{\rm PBH}=D/\sqrt{n_{\rm Bm}}$. Hence, under the simplifying assumption that all BH mergers eject the same amount of morsels of the same mass, one can estimate the limit on the BH morsel fraction by comparing the rate of BH mergers measured by LIGO/VIRGO [61], $\rho_{\rm LV}=24^{+14}_{-9}\;\mbox{Gpc}^{-3}\ \mbox{yr}^{-1}$, to the rescaled limit on PBH densities from HAWC, $\rho_{\rm PBH}=3400\;\mbox{pc}^{-3}\ \mbox{yr}^{-1}$. The estimated lower limit on the total morsel mass is shown in Fig. 2 in green, where the band comes from the uncertainty in the BH merger rate. A dedicated search, taking into account the morsel mass distributions for different merger events, can provide a more reliable limit.
@@
Neutrino observatories like ANTARES and IceCube have also monitored BH mergers in search for neutrino counterparts [62]. Both ANTARES [63, 64] and IceCube [65] search for neutrino events within a window of ±$\pm 500$ sec around the BH merger, finding no excess over the expected backgrounds. However, the limits in terms of fluxes are orders of magnitude above the ones for photons, requiring typical source luminosity of $10^{51}$ erg sec-1. As the BH morsels predict similar fluxes in photons and neutrinos, searches for the latter remain not competitive.
2.2 Uncertainties and Sensitivity Analysis
The predictions presented in the previous section rely on numerical computations from BlackHawk and on assumptions regarding the distribution, number, and initial conditions of the BH morsels. Here, we recap the uncertainties stemming from:
  • BlackHawk Limitations: We note that the BlackHawk software used for our calculations relies on semiclassical assumptions and a fixed (non-backreacting) background. Hence, it does not capture possible quantum gravity or dynamical spacetime effects [30, 31]. Another source of uncertainty stems from the hadronisation model, which typically affects the results at the $20$ level.
  • Morsel Mass Distribution: We considered a single mass distribution, with typical morsel masses ranging from $10^{7}$ to $10^{9}$ kg in order to have an observable time-delay from the merger. Changing the initial mass distribution would alter the predicted flux normalization and duration.
  • Spin Effects: While we assume non-rotating morsels, including initial spin decreases the Hawking temperature, shifts the spectrum to lower frequencies, and introduces anisotropic emission. Furthermore, the evaporation is slowed down, however the spin dissipation is faster [33]. Hence, including spin effects should lead to minor modifications of our results.
  • Merger Rate Uncertainties: Published LIGO/VIRGO rates vary by an order of magnitude [61], directly affecting the expected detection rate for non-coincidence searches. This uncertainty has been included in the green curve in Fig. 2.
As a first estimate, we indicate the impact of parameter variation in Table 1. Future work should include full Monte Carlo parameter scans and systematic error propagation.
Parameter Variation Flux Impact (approx.)
Numerical (BlackHawk) hadronisation $\sim 20$
Morsel mass distr. $10^{7}$$10^{9}$ kg Duration: min–yr; Flux amplitude: $\sim$10$\times$
[…]
Table 1: Estimated sensitivity of predicted photon flux to key model parameters.
2.3 Supermassive BH mergers